### v1

# references.bib

Abstract
Integrated Sensing and Communication (ISAC) systems are prone to privacy violations, once they aim at handling sensitive identifiable information in several applications. This paper raises the necessity of implementing privacy-preservation measures on the design of cell-free massive multiple-input multiple-output ISAC systems. To that purpose, given an adversary model, we propose an iterative framework of two blocks, precoder design and access point selection. The precoder design aims at maximizing the signal-to-interference-plus-noise ratio at the sensing receivers given communication constraints. The access point selection aims at minimizing the mutual information between the received signal at users and the sensing signal, by rearranging the access points that transmit ISAC-signals and the sensing receivers. Results show that a reduction in the probability of detection by the adversary is obtained with this method.
Index Terms:
ISAC, cell-free MIMO, Privacy
I Introduction
The evolution of multiple input multiple output (MIMO) technologies toward distributed implementations, in a cell-free setting, will serve as a infrastructure to support distributed integrated sensing and communication (ISAC) systems and networked sensing implementations [behdad-multi-static-2023]. While a significant progress has been done in the literature to propose the best strategies for conciliating communication and sensing requirements, a primary concern in the design of the physical layer of ISAC is security and privacy. Particularly, the ability of malicious parties to use sensing information to track and locate targets can lead to the emerging of unprecedented attacks.
While eavesdropping attacks have been addressed in a number of physical layer security-based solutions [9737364], privacy attacks at the physical layer and countermeasures remain barely explored. A privacy attack model is proposed in [da-silva-multi-static-2023], which consider an internal attacker that leverages the signals received from ISAC access points (APs) to estimate the position of targets by inferring the directions the APs originally beamforming their sensing signals toward these targets. Moreover, the work in [10587082] proposes a secure sensing scheme to prevent anauthorized sensing from a malicious communication user via beamforming design.
Therefore, inspired by the attack model in [da-silva-multi-static-2023], this paper proposes an iterative framework of two blocks, precoder design and access point selection. The precoder design aims at maximizing the signal-to-interference-plus-noise ratio (SINR) at the sensing receivers given communication constraints. The access point selection aims at minimizing the mutual information (MI) between the received signal at users and the sensing signal, by rearranging the access points that transmit ISAC signals and the sensing receivers.
II System Model
Consider a centralized cell-free massive MIMO system of $N_{\text{Tx}}$ transmitting ISAC APs, with $M$ isotropic antennas, jointly serving $N_{\text{UE}}$ $K$-antenna users. However, while the user terminals may consist of $K$ antennas, a decoder designed by a centralized unit would attempt to extract the data only meant for the user which is not desirable by the adversarial model described later. Therefore, a decoder design can be disregarded and the multiple antenna users can be modelled as single antenna users receiving additional independent observations of the signal. Additionally, $N_{\text{Rx}}$ $M$-antenna receiver APs are performing sensing of a static target present in the system, which is illuminated by the ISAC transmitters. This system is depicted in 1.
Figure 1: System Model
The received signal, $y_{i}[n]$, at user $i=1,2,...,N_{\text{UE}}$ with communication symbols $s_{i}[n]$, where $n=1,2...,N$ denotes the $n$th sample is
$\begin{split}y_{i}[n]=\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,i}s_{i}[n]}_{\text{Desired Signal}}+\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,k}s_{k}[n]}_{\text{Communication Interference}}\\ +\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,t}s_{s}[n]}_{\text{Sensing Interference}}+\underbrace{n_{i}[n]}_{\text{Noise}},\end{split}$ (1)
where $\mathbf{h}_{j,i}$ and $\mathbf{w}_{j,i}$ are $M\times 1$ column vectors for the channel coefficients and the precoder, from AP $j$ to user $i$, respectively. Similarly, $\mathbf{w}_{j,t}$ is the sensing precoding vector from AP $j$ to the target $t$. Since the channel is assumed to be Rayleigh fading, the channel coefficients $\mathbf{h}_{j,i}\in\mathbb{C}^{M}$ consists of independent and identically distributed $\mathcal{CN}(0,\beta_{j,i})$ entries, where $\beta_{j,i}$ is the channel gain between AP $j$ and user $i$. Also, $n_{i}[n]\sim\mathcal{CN}(0,\sigma^{2}_{n})$ represents the received additive white Gaussian noise (AWGN) at user $i$, and $s_{s}[n]$ represents the sensing probing signal.
We can write the transmitted $M\times 1$ signals $\mathbf{x}_{j}[n]$ at each AP $j=1,2,...,N_{\text{Tx}}$ as
$\mathbf{x}_{j}[n]=\sum\limits_{i=1}^{N_{\text{UE}}}\mathbf{w}_{j,i}s_{i}[n]+\mathbf{w}_{j,t}s_{s}[n].$ (2)
Further, an expression for the received signal at the receiver AP is formulated, which depends on the antenna array response vector of both the transmitter and the receiver. The $M\times 1$ antenna array response vector $\mathbf{a}(\theta)$ for a uniform linear array is defined as $\mathbf{a}(\theta)=$$\left[1\quad e^{j\pi\cos{\theta}}\quad e^{j2\pi\cos{\theta}}\quad\dots\quad e^{j(M-1)\pi\cos{\theta}}\right]^{T},$ where $\theta$ is the horizontal angle of arrival/departure of the signals. The received $M\times 1$ signal vector $\mathbf{y}[n]$ at receiver AP $r$ can be written as
$\begin{split}\mathbf{y}_{r}[n]=\sum\limits_{j=1}^{N_{\text{Tx}}}\alpha_{j,r}\sqrt{\beta_{j,r}}\mathbf{a}(\theta_{r})\mathbf{a}(\theta_{j})^{T}\mathbf{x}_{j}[n]\\ +\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{H}_{j,r}\mathbf{x}_{j}[n]+\mathbf{n}[n].\end{split}$ (3)
In (3) $\alpha_{j,r}\in\mathbb{C}$ is the radar cross section of the target following the Swerling-I model and $\beta_{j,r}$ is the channel gain between AP $j$, the target, and the receiver AP, with the path’s phase shift being included in $\alpha_{j,r}$. Also, $\mathbf{n}[n]\in\mathbb{C}^{M}$ is the AWGN vector with independent and identically distributed $\mathcal{CN}(0,\sigma_{n}^{2})$ entries. The angles $\theta_{j}$ and $\theta_{r}$ are the angles of departure from transmitter AP $j$, and the reception angle of the echoes from the target to the receiver AP, respectively. $\mathbf{H}_{j,r}\in\mathbb{C}^{M\times M}$ are the channel coefficients for the unknown non-line-of-sight paths caused by temporary objects.
III Adversary Model
The adversary is assumed to be one of the communication users that maliciously estimates the position of the target by inferring the directions of the beampatterns from the transmitting APs. Different from [da-silva-multi-static-2023], herein we propose a gradient descent method to minimize the mean square error of the distance from a line projected from each known AP location, in the direction of the estimated beampattern. Then, if the target is within a predetermined radius, $R_{\text{GD}}$, from the estimated location, the estimate is deemed successful in determining the target position. Mathematically, the process can be described as
$\mathbf{q}_{p+1}=\mathbf{q}_{p}-\eta\nabla\text{MSE}(\mathbf{q}_{p}),$ (4)
$\text{MSE}(\mathbf{q})=\dfrac{1}{N_{l}}\sum\limits_{i=1}^{N_{l}}||\mathbf{q}-\mathbf{y}_{i}||^{2}$ (5)
where $\mathbf{q}$ is the estimated point, $p$ is the iteration index, $\eta$ and $\nabla$ denotes the step size and gradient operator, respectively, $N_{l}$ is the number of lines, and $\mathbf{y}_{i}$ is the closest point of each line to the estimated point.
Since the peak of each AP’s beampattern is its sensing signal angle, the adversary tries to replicate this beampattern. This can be done by taking advantage of the precoder being a hidden variable within the transmitted signal $\mathbf{x}_{j}$, as seen in (2), and can be studied in [10587082]. Therefore, the outer product of the transmitted signals, i.e, $\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}$, will contain information about the precoder, which in turns allows the adversary to sweep $\mathbf{a}(\theta_{s})^{H}\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\mathbf{a}(\theta_{s})$ over an angular search area, $\theta_{s}$, to estimate the beampattern and finding its peak. Since the locations of the APs are known by the adversary, $\theta_{s}$ is therefore also assumed to be, at least, within the same half-plane as the correct angle.
For estimating the transmitted signal, the expectation maximization (EM) algorithm [moon-expectation-maximization-1996] is used. In this case, the log-likelihood function used in the EM procedure is $\log(p(\widetilde{y}_{a,j};\mathbf{x}_{j}))$, where $p(\widetilde{y}_{a,j};\mathbf{x}_{j})$ is the probability density function of the received signal $\widetilde{y}_{a,j}$ at the adversary $a$, from AP $j$, given the transmitted signal $\mathbf{x}_{j}$.
The EM approach is employed due to considering the actual channel between the adversary and AP $j$, as a latent variable within the likelihood function. For convenience, the channel estimate used in the initial iteration of the EM algorithm is modelled to be the actual channel, with added Gaussian noise to mimic some channel estimation error. Thus, the channel estimate $\widehat{\mathbf{H}}_{a,j}$ is a Gaussian random variable with mean $\mathbf{H}_{a,j}$ and variance $\sigma^{2}_{H}\mathbf{I}$, where $\sigma^{2}_{H}$ is the channel estimation error variance.
Following [da-silva-multi-static-2023] and [zhang-deep-2022], $u$ is introduced as an additional latent variable, allowing the definitions $\mathbf{h}_{a,j}=\{\mathbf{H}_{a,j},u\}$, as well as $\boldsymbol{\theta}=\{\mathbf{x}_{j},\nu\}$ to be made, with $\nu$ being the degree of freedom of an underlying gamma distribution (details are in [zhang-deep-2022]). Thereafter, by establishing $q(\mathbf{h}_{a,j})$ as a variational arbitrary distribution, the maximizable expression can be defined as
[…]
III-1 Expectation Step
The expectation step maximizes $\mathcal{F}(\widetilde{y}_{a,j};q,\boldsymbol{\theta})$, with respect to the latent variable, which in this case is $q(\mathbf{h}_{a,j})$. Further, through studying [da-silva-multi-static-2023], $\mathcal{F}(\widetilde{y}_{a,j};q,\boldsymbol{\theta})$ is maximized in this step by setting
$\mathbb{E}_{q}[\mathbf{H}_{a,j}]=\left[\boldsymbol{\Omega}_{\mathbf{H}}\left(\frac{1}{\sigma_{H}^{2}}\widehat{\mathbf{H}}_{a,j}^{T}+\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\frac{\mathbb{E}_{q}[u]}{\sigma_{n}^{2}}\right)\right]^{H},$ (7)
$\boldsymbol{\Omega}_{\mathbf{H}}=\left(\frac{1}{\sigma_{H}^{2}}\mathbf{I}+\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\frac{\mathbb{E}_{q}[u]}{\sigma^{2}_{n}}\right)^{-1},$ (8)
$\mathbb{E}_{q}[u]=\mathbb{E}\left[\mathcal{G}\left(\frac{v+2M_{r}}{2},\frac{v+C}{2}\right)\right],$ (9)
where $M_{r}$ and $M$ is the number of antennas at the receiver and transmitter, respectively. Also, $\mathbb{E}_{q}$ denotes the expected value of the variational distribution $q$ of the variables. Finally, $C$ in (9) is defined as
[…]
It can be noted that there is a dependency between $\mathbb{E}_{q}[u]$ and $\mathbb{E}_{q}[\mathbf{H}_{a,j}]$, which is mitigated by using the previous iteration’s $\mathbb{E}_{q}[u]$ in the calculations, and in the case of it being the first iteration, it is randomly initiated.
III-2 Maximization Step
This step involves determining the estimate of $\mathbf{x}_{j}$ by maximizing $\mathcal{F}(\widetilde{y}_{a,j};q,\boldsymbol{\theta})$ with regards to $\boldsymbol{\theta}$, and it is shown in [zhang-deep-2022] that the transmitted signal from AP $j$, may be determined using
$\begin{split}\widehat{\mathbf{x}}_{j}=\operatorname*{argmin}_{\textbf{x}_{j}}\left[(\widetilde{y}_{a,j}-\mathbb{E}_{q}[\mathbf{H}_{a,j}]\textbf{x}_{j})^{\,H}(\widetilde{y}_{a,j}-\mathbb{E}_{q}[\mathbf{H}_{a,j}]\textbf{x}_{j})\right.\\ +\left.\textbf{x}_{j}^{\,H}\boldsymbol{\Omega}_{\mathbf{H}}M\textbf{x}_{j}\right].\end{split}$ (11)
These steps are then repeated until the difference of two consecutive iterations of the estimate $\widehat{\mathbf{x}}_{j}$ are sufficiently small, or when the predetermined maximum number of iterations has been achieved.
IV Privacy-Preserving Framework
For a cell-free MIMO ISAC system, we propose a joint, two block, precoder and AP selection framework for privacy-preservation, which utilizes alternating optimization of the two blocks. For a given receiver/transmitter AP configuration, the first block finds the precoding vectors that maximize the SINR at the sensing receivers. The second block minimizes the probability of detection of the target from the adversary, by altering the AP configuration. Henceforth, this probability is denoted $P_{D}$. As a convergence criteria, the procedure stops when two iterations leave the AP configuration unchanged. The first iteration assumes all APs as transmitters.
IV-A Precoder Design
Given a set of APs with predetermined receiver/transmitter duties, the problem of maximization of the received SINR at sensing APs can be formulated as
$\max_{\mathbf{W}}\quad$ $\gamma_{s}(\mathbf{W})$ (12a)
s.t. $\gamma_{\text{UE}_{i}}\geq\gamma,\forall i$ (12b)
$P_{j,a}\leq P_{max},\forall j,\forall a,$ (12c)
where $\mathbf{W}$ is a $MN_{\text{Tx}}\times(N_{\text{UE}}+1)$ matrix consisting of each transmitter APs $M\times(N_{\text{UE}}+1)$ precoder matrix, denoted $\mathbf{W}_{j}$ for AP $j=1,2,...,N_{\text{Tx}}$. In turn, $\mathbf{W}_{j}$ consists of AP $j$’s concatenated precoder vectors for each user plus the sensing precoder, that is $\mathbf{W}_{j}=\left[\mathbf{w}_{j,1},\mathbf{w}_{j,2},...,\mathbf{w}_{j,N_{\text{UE}}},\mathbf{w}_{j,t}\right]$. Hence, the aggregate precoder matrix $\mathbf{W}=\left[\mathbf{W}_{1}^{T},\mathbf{W}_{2}^{T},...,\mathbf{W}_{N_{\text{Tx}}}^{T}\right]^{T}$ columns represent the total precoding vector for a user from every transmitter AP, where the last column represents the target. Further, $\gamma_{\text{UE}_{i}}$ and $\gamma$ represents the total received SINR per user and the minimum SINR threshold for a user, respectively. $P_{j,a}$ is the transmit power from AP $j$’s antenna $a$, and $P_{max}$ is the maximum allowed transmission power for each antenna. Therefore, $P_{j,a}$ can be computed by taking the norm square of each row of $\mathbf{W}$.
Similar to [behdad-multi-static-2023], by taking the norm square of the channel gain of the desired signal in (3), $\gamma_{s}(\mathbf{W})$ may be expanded as
$\gamma_{s}(\mathbf{W})=\frac{\displaystyle\sum\limits_{r=1}^{N_{\text{Rx}}}\sum\limits_{n=1}^{N}\mathbf{s}[n]^{H}\mathbf{W}^{H}\mathbf{A}_{r}\mathbf{W}\mathbf{s}[n]}{\displaystyle N_{\text{Rx}}NM\sigma^{2}_{n}}.$ (13)
by using the the aggregated $MN_{\text{Tx}}\times MN_{\text{Tx}}$ matrix for each receiver $r$
$\mathbf{A}_{r}=\begin{bmatrix}\mathbf{A}_{1,1}&\dots&\mathbf{A}_{1,N_{\text{Tx}}}\\ \vdots&\ddots&\vdots\\ \mathbf{A}_{N_{\text{Tx}},1}&\dots&\mathbf{A}_{N_{\text{Tx}},N_{\text{Tx}}}\end{bmatrix},$
with $\mathbf{A}_{j,m}$ being the $M\times M$ matrices of transmitter AP $j$ and $m$ defined by
$\mathbf{A}_{j,m}=\sqrt{\beta_{j,r}\beta_{m,r}}\mathbf{a}(\theta_{j})^{*}\mathbf{a}(\theta_{r})^{H}\alpha_{j,r}^{*}\alpha_{m,r}\mathbf{a}(\theta_{r})\mathbf{a}(\theta_{m})^{T}.$
In (13), $N$ denotes the number of symbols that are being transmitted to each user, and, consequently, the number of samples in the sensing probing signal. Moreover, $\mathbf{s}[n]$ is the vector of transmitted symbols for all users and the sensing signal, i.e $\left[s_{1}[n],...,s_{N_{\text{UE}}}[n],s_{s}[n]\right]$.
Since the aim is to optimize $\gamma_{s}$ with respect to $\mathbf{W}$, the Concave-Convex Procedure (CCP) is considered and thus, the objective function must be concave. For this purpose the first order Taylor approximation is applied to the term $\mathbf{W}^{H}\mathbf{A}_{r}\mathbf{W}$ around the result of the previous iteration of the procedure, resulting in
$P_{1}^{r}(\mathbf{W}_{p})=2\mathbf{W}^{H}_{p-1}\mathbf{A}_{r}\left(\mathbf{W}_{p}-\mathbf{W}_{p-1}\right)+\mathbf{W}^{H}_{p-1}\mathbf{A}_{r}\mathbf{W}_{p-1}.$ (14)
and thus a final expression for the objective function is obtained as
$\gamma_{s}(\mathbf{W})=\frac{\displaystyle\sum\limits_{r=1}^{N_{\text{Rx}}}\sum\limits_{n=1}^{N}\mathbf{s}[n]^{H}P_{1}^{r}(\mathbf{W})\mathbf{s}[n]}{\displaystyle N_{\text{Rx}}NM\sigma^{2}_{n}}.$ (15)
In (14), $p$ refers to the iteration of the CCP, hence $p-1$ is the $\mathbf{W}$ from the previous iteration and a random matrix is generated for $p=0$.
A similar approach is used to transform the communication SINR constraint into a convex function where the function $\gamma_{\text{UE}_{i}}$ is expanded to
$\gamma_{\text{UE}_{i}}=\frac{\displaystyle\left|\mathbf{h}_{i}^{H}\mathbf{w}_{i}\right|^{2}}{\displaystyle\left|\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{i}^{H}\mathbf{w}_{k}\right|^{2}+\left|\mathbf{h}_{i}^{H}\mathbf{w}_{t}\right|^{2}+\sigma^{2}_{n}},$ (16)
by using the substitutions $\mathbf{h}_{i}=\left[\mathbf{h}^{T}_{1,i},\mathbf{h}^{T}_{2,i},...,\mathbf{h}^{T}_{N_{\text{Tx}},i}\right]^{T}$ and
$\mathbf{w}_{i}=\left[\mathbf{w}^{T}_{1,i},\mathbf{w}^{T}_{2,i},...,\mathbf{w}^{T}_{N_{\text{Tx}},i}\right]^{T}$.
Again the first order Taylor approximation is applied around the previous iteration’s value of $\mathbf{w}_{i}$. Also, the slack variables $\tau_{d}$ and $\tau_{n}$ for the denominator and numerator in (16), respectively, are introduced. The final expression for the constraint in (12b) is therefore
$(\mathbf{w}_{i}^{H})_{(p-1)}\mathbf{h}_{i}\mathbf{h}_{i}^{H}(\mathbf{w}_{i})_{(p-1)}\\ +2(\mathbf{w}_{i}^{H})_{(p-1)}\mathbf{h}_{i}\mathbf{h}_{i}^{H}((\mathbf{w}_{i})_{p}-(\mathbf{w}_{i})_{(p-1)})\geq\tau_{n}$ (17a)
$\left|\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{i}^{H}\mathbf{w}_{k}\right|^{2}+\left|\mathbf{h}_{i}^{H}\mathbf{w}_{t}\right|^{2}+\sigma^{2}_{n}\leq\tau_{d}$ (17b)
$\tau_{n}\geq\tau_{d}\gamma.$ (17c)
IV-B Transmitter and Receiver Access Point Selection
The proposed method to minimize $P_{D}$ involves a minimization of the MI w.r.t. APs being ISAC-signal receivers or transmitters.
Consider the set of all APs (initially considered as transmitters), $\mathcal{A}=$$\{a_{l}\}_{1}^{N_{\text{AP}}}$ where $a_{l}$ is each AP $l=1,2,...,N_{AP}$, and $N_{AP}$ is the total number of APs. The problem formulation then becomes
$\min_{\mathcal{R}\subset\mathcal{A}}\quad\sum\limits_{i=1}^{N_{\text{UE}}}\sum\limits_{l=1}^{N_{\text{AP}}}I(y_{i};\mathbf{x}^{s}_{l}),$ (18)
where $\mathcal{R}$ is the set of sensing receiver APs.
Following [marzetta-fundamentals-2016], the MI can be upper bounded by
$I(y_{i};\mathbf{x}^{s}_{l})\leq\log_{2}\left[\det\left(\mathbf{I}+\mathbf{H}_{i,l}\mathbf{x}^{s}_{l}(\mathbf{x}^{s}_{l})^{H}\mathbf{H}_{i,l}^{H}\right)\right],$ (19)
Finally, by defining $\mathbf{l}_{\text{Rx}}=\left[l_{1},l_{2},...,l_{N_{\text{Rx}}}\right]^{T}$ as the unique indexes of the receiver APs in $\mathcal{A}$, the solution to (18) becomes
=$\mathcal{R}=\left\{a_{l}\in\mathcal{A}\,\,\forall l\in\mathbf{l}_{\text{Rx}}:\operatorname*{argmax}_{\mathbf{l}_{\text{Rx}}}\sum\limits_{l\in\mathbf{l}_{\text{Rx}}}\sum\limits_{i=1}^{N_{\text{UE}}}I(y_{i};\mathbf{x}^{s}_{l})\right\}$ (20)
V Results
Results presented in this section are produced using the parameters in Table I unless specified. The system is randomized, with certain constraints [^1] https://gitlab.liu.se/henak56/privacy-for-isac, for each of the $500$ trials used for each data point in all plots.
TABLE I: Simulation parameters.
Parameter Value Parameter Value
$N$ $16$ $N_{\text{UE}}$ $5$
Modulation Order $16$ $K$ $4$
Path Loss Exponent $2.8$ $P_{max}$ $40$ dBm
Max EM Iterations $100$ $\gamma$ $3$ dB
Max CCP Iterations $100$ $\sigma_{H}$ $-40$ dB
EM Convergence Threshold $1\cdot 10^{-10}$ $\alpha_{j,r}$ $10$ dB
CCP Convergence Threshold $0.1$ $\sigma_{n}$ $-94$ dBm
$N_{\text{AP}}$ $5$ $\eta$ $1$
$M$ $64$ $R_{\text{GD}}$ $10$
Figure 2: Probability of target detection, $P_{D}$, with varying min power constraint.
Figure 3: Probability of detection, $P_{D}$, for varying number of receiver APs.
Figure 2 shows $P_{D}$ when using no MI minimization, versus using the framework proposed in this paper. As expected, $P_{D}$ is greater when more power is available to the system, due to the sensing beams being more prevalent within the transmitter beampatterns. Moreover, when the framework is limited to optimizing w.r.t. only one receiver AP, a ceiling in reduction of $P_{D}$ seems to be reached, independent of alterations of the maximum power usage constraint.
As seen in 3, $P_{D}$ decreases with more receiver APs, as there are less ISAC transmitters to triangulate an estimated target position. Furthermore, the difference between the two cases increases as more APs are acting as receivers, which demonstrates a privacy-preservation gain from having more degrees of freedom with the developed method. It is worthwhile to mention that, in both cases, the communication SINR and maximum power consumption constraints are met.
VI Conclusions
Attempting to reduce the probability of detection of a target within a ISAC system, from an adversarial user of the system, was the aim of this paper. Using the proposed MI minimization technique, the detection probability stayed the same or decreased in all cases, but further investigations on the trade-offs between computational efforts and allowing for lower sensing SINR are still to be carried out.
[…]