### v2

# references.bib

Abstract
Integrated Sensing and Communication (ISAC) systems are prone to privacy violations, once they aim at handling sensitive identifiable information in several applications. This paper raises the necessity of implementing privacy-preservation measures on the design of cell-free massive multiple-input multiple-output ISAC systems. To that purpose, given an adversary model, we propose an iterative framework of two blocks, precoder design and access point selection. The precoder design aims at maximizing the signal-to-interference-plus-noise ratio at the sensing receivers given communication constraints. The access point selection aims at minimizing the mutual information between the received signal at users and the sensing signal, by rearranging the access points that transmit ISAC-signals and the sensing receivers. Results show that a reduction in the probability of detection by the adversary is obtained with this method.
Index Terms:
ISAC, cell-free MIMO, Privacy
I Introduction
The evolution of multiple input multiple output (MIMO) technologies toward distributed implementations, in a cell-free setting, will serve as a infrastructure to support distributed integrated sensing and communication (ISAC) systems and networked sensing implementations [behdad-multi-static-2023]. While significant progress has been done in the literature to propose the best strategies for conciliating communication and sensing requirements, a primary concern in the design of the physical layer of ISAC is security and privacy [qu_privacy_2024]. Particularly, the ability of malicious parties to use sensing information to track and locate targets can lead to the emerging of unprecedented attacks [wang_anti-malicious_2024]. However, methods for combating the security and privacy concerns inherent in ISAC has been proposed. For example [su_secure_2023, 10587082], where the sensing functionality is utilized to locate attackers, and hamper eavesdroppers by artificial noise transmission. While eavesdropping attacks have been addressed in a number of physical layer security-based solutions [9737364], privacy attacks at the physical layer and countermeasures remain barely explored. A privacy attack model is proposed in [da-silva-multi-static-2023], which consider an internal attacker that leverages the signals received from ISAC access points (APs) to estimate the position of targets by inferring the directions the APs originally beamforming their sensing signals toward these targets. Moreover, the work in [10587082] proposes a secure sensing scheme to prevent unauthorized sensing from a malicious communication user via beamforming design.
Note that, different from [da-silva-multi-static-2023], we propose a different version of the attack model. Additionally, we propose an iterative framework of two blocks, precoder design and access point selection, to reduce the impact of the attack. The precoder design aims at maximizing the signal-to-interference-plus-noise ratio (SINR) at the sensing receivers given communication constraints. The access point selection aims at minimizing the mutual information (MI) between the received signal at users and the sensing signal, by rearranging the access points that transmit ISAC signals and the sensing receivers.
II System Model
Consider a centralized cell-free massive MIMO system of $N_{\text{Tx}}$ transmitting ISAC APs, with $M$ isotropic antennas, jointly serving $N_{\text{UE}}$ $K$-antenna users. Decoders designed by the centralized control unit would aim to extract data meant for the user, hence, a decoder design can be disregarded and multiple antennas at the users can be modeled as receiving additional independent observations of the signal. Additionally, $N_{\text{Rx}}$ $M$-antenna receiver APs are performing sensing of a static target present in the system, which is illuminated by the ISAC transmitters. This system is depicted in Figure 1.
Figure 1: System Model
The received signal, $y_{i}[n]$, at user $i=1,2,...,N_{\text{UE}}$ with communication symbols $s_{i}[n]$, where $n=1,2...,N$ denotes the $n$th sample is
$y_{i}[n]=\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,i}s_{i}[n]}_{\text{Desired Signal}}+\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,k}s_{k}[n]}_{\text{Communication Interference}}\\ +\underbrace{\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{h}_{j,i}^{H}\mathbf{w}_{j,t}s_{s}[n]}_{\text{Sensing Interference}}+\underbrace{n_{i}[n]}_{\text{Noise}},$ (1)
where $\mathbf{h}_{j,i}$ and $\mathbf{w}_{j,i}$ are $M\times 1$ column vectors for the channel coefficients and the precoder, from AP $j$ to user $i$, respectively. Similarly, $\mathbf{w}_{j,t}$ is the sensing precoding vector from AP $j$ to the target $t$. Since the channel is assumed to be Rayleigh fading, the channel coefficients $\mathbf{h}_{j,i}\in\mathbb{C}^{M}$ consists of independent and identically distributed $\mathcal{CN}(0,\beta_{j,i})$ entries, where $\beta_{j,i}$ is the channel gain between AP $j$ and user $i$. Also, $n_{i}[n]\sim\mathcal{CN}(0,\sigma^{2}_{n})$ represents the received additive white Gaussian noise (AWGN) at user $i$, and $s_{s}[n]$ represents the sensing probing signal.
We can write the transmitted $M\times 1$ signals $\mathbf{x}_{j}[n]$ at each AP $j=1,2,...,N_{\text{Tx}}$ as
$\mathbf{x}_{j}[n]=\sum\limits_{i=1}^{N_{\text{UE}}}\mathbf{w}_{j,i}s_{i}[n]+\mathbf{w}_{j,t}s_{s}[n].$ (2)
Further, an expression for the received signal at the receiver AP is formulated, which depends on the antenna array response vector of both the transmitter and the receiver. The $M\times 1$ antenna array response vector $\mathbf{a}(\theta)$ for a uniform linear array is defined as $\mathbf{a}(\theta)=$$\left[1\quad e^{j\pi\cos{\theta}}\quad e^{j2\pi\cos{\theta}}\quad\dots\quad e^{j(M-1)\pi\cos{\theta}}\right]^{T},$ where $\theta$ is the horizontal angle of arrival/departure of the signals. The received $M\times 1$ signal vector $\mathbf{y}[n]$ at receiver AP $r$ can be written as
$\mathbf{y}_{r}[n]=\sum\limits_{j=1}^{N_{\text{Tx}}}\alpha_{j,r}\sqrt{\beta_{j,r}}\mathbf{a}(\theta_{r})\mathbf{a}(\theta_{j})^{T}\mathbf{x}_{j}[n]\\ +\sum\limits_{j=1}^{N_{\text{Tx}}}\mathbf{H}_{j,r}\mathbf{x}_{j}[n]+\mathbf{n}[n].$ (3)
In (3) $\alpha_{j,r}\in\mathbb{C}$ is the radar cross section of the target following the Swerling-I model and $\beta_{j,r}$ is the channel gain between AP $j$, the target, and the receiver AP, with the path’s phase shift being included in $\alpha_{j,r}$. Also, $\mathbf{n}[n]\in\mathbb{C}^{M}$ is the AWGN vector with independent and identically distributed $\mathcal{CN}(0,\sigma_{n}^{2})$ entries. The angles $\theta_{j}$ and $\theta_{r}$ are the angles of departure from transmitter AP $j$, and the reception angle of the echoes from the target to the receiver AP, respectively. $\mathbf{H}_{j,r}\in\mathbb{C}^{M\times M}$ are the channel coefficients for the unknown non-line-of-sight paths caused by temporary objects.
III Adversary Model
The adversary is assumed to be one of the communication users that maliciously estimates the position of the target by inferring the directions of the beampatterns from the transmitting APs. Henceforth, for the adversary model, the sample index $n$ is dropped and it is implied that the following methods are repeated for every sample.
Different from [da-silva-multi-static-2023], herein we propose a gradient descent method to minimize the mean square error of the distance from a line projected from each known AP location, in the direction of the estimated beampattern. Then, if the target is within a predetermined radius, $R_{\text{GD}}$, from the estimated location, the estimate is deemed successful in determining the target position. Mathematically, the process can be described as
$\mathbf{q}_{p+1}=\mathbf{q}_{p}-\eta\nabla\text{MSE}(\mathbf{q}_{p}),$ (4)
$\text{MSE}(\mathbf{q})=\dfrac{1}{N_{l}}\sum\limits_{i=1}^{N_{l}}||\mathbf{q}-\mathbf{y}_{i}||^{2},$ (5)
where $\mathbf{q}$ is the estimated point, $p$ is the iteration index, $\eta$ and $\nabla$ denotes the step size and gradient operator, respectively, $N_{l}$ is the number of lines, and $\mathbf{y}_{i}$ is the closest point of each line to the estimated point.
Since the peak of each AP’s beampattern is its sensing signal angle, the adversary tries to replicate this beampattern. This can be done by taking advantage of the precoder being a hidden variable within the transmitted signal $\mathbf{x}_{j}$, as seen in (2), and can be studied in [10587082]. Therefore, the outer product of the transmitted signals, i.e, $\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}$, will contain information about the precoder, which in turns allows the adversary to sweep $\mathbf{a}(\theta_{s})^{H}\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\mathbf{a}(\theta_{s})$ over an angular search area, $\theta_{s}$, to estimate the beampattern and finding its peak [liu_joint_2020]. Since the locations of the APs are known by the adversary, $\theta_{s}$ is therefore also assumed to be, at least, within the same half-plane as the correct angle.
For estimating the transmitted signal, the expectation maximization (EM) algorithm [moon-expectation-maximization-1996] is used. In this case, the log-likelihood function used in the EM procedure is $\log(p(y_{j,a};\mathbf{x}_{j}))$, where $p(y_{j,a};\mathbf{x}_{j})$ is the probability density function of the received signal $y_{j,a}$ at the adversary $a$, from AP $j$, given the transmitted signal $\mathbf{x}_{j}$.
The EM approach is employed due to considering the actual channel, $\mathbf{h}_{j,a}\in\mathbb{C}^{M}$, between the adversary and AP $j$, as a latent variable within the likelihood function. For convenience, the channel estimate used in the initial iteration of the EM algorithm is modelled to be the actual channel, with added Gaussian noise to mimic some channel estimation error. Thus, the channel estimate $\widehat{\mathbf{h}}_{j,a}$ is a Gaussian random variable with mean $\mathbf{h}_{j,a}$ and covariance matrix $\sigma^{2}_{h}\mathbf{I}_{M}$, where $\sigma^{2}_{h}$ is the channel estimation error variance, and $\mathbf{I}_{M}$ is an identity matrix of size $M\times M$.
Following [da-silva-multi-static-2023] and [zhang-deep-2022], $u$ is introduced as an additional latent variable, allowing the definitions $\mathbf{z}_{j,a}=\{\mathbf{h}_{j,a},u\}$, as well as $\boldsymbol{\theta}=\{\mathbf{x}_{j},\nu\}$ to be made, with $\nu$ being the degree of freedom of an underlying Student’s $t$-distribution (details are in [zhang-deep-2022]). Thereafter, by establishing $q(\mathbf{z}_{j,a})$ as a variational arbitrary distribution, the maximizable expression can be defined as
[…]
III-1 Expectation Step
The expectation step maximizes $\mathcal{F}(y_{j,a};q,\boldsymbol{\theta})$, with respect to the latent variable, which in this case is $q(\mathbf{z}_{j,a})$. Further, through studying [da-silva-multi-static-2023], $\mathcal{F}(y_{j,a};q,\boldsymbol{\theta})$ is maximized in this step by setting
$\mathbb{E}_{q}[\mathbf{h}_{j,a}]=\left[\boldsymbol{\Omega}_{\mathbf{h}}\left(\frac{1}{\sigma_{h}^{2}}\widehat{\mathbf{h}}_{j,a}+\mathbf{x}_{j}y_{j,a}\frac{\mathbb{E}_{q}[u]}{\sigma_{n}^{2}}\right)\right],$ (7)
$\boldsymbol{\Omega}_{\mathbf{h}}=\left(\frac{1}{\sigma_{h}^{2}}\mathbf{I}_{M}+\mathbf{x}_{j}\mathbf{x}_{j}^{\,H}\frac{\mathbb{E}_{q}[u]}{\sigma^{2}_{n}}\right)^{-1},$ (8)
$\mathbb{E}_{q}[u]=\mathbb{E}\left[\mathcal{G}\left(\frac{\nu+2K}{2},\frac{\nu+C}{2}\right)\right],$ (9)
where $\mathbb{E}_{q}$ denotes the expected value of the variational distribution $q$ of the variables, and $\mathbf{I}_{M}$ is an $M\times M$ identity matrix. Finally, $C$ in (9) is defined as
$C=\frac{1}{\sigma^{2}_{n}}\left[\left(y_{j,a}-\mathbb{E}_{q}[\mathbf{h}_{j,a}]^{H}\mathbf{x}_{j}\right)^{\!2}\,+M\mathbf{x}_{j}^{H}\boldsymbol{\Omega}_{\mathbf{h}}\mathbf{x}_{j}\right].$ (10)
It can be noted that there is a dependency between $\mathbb{E}_{q}[u]$ and $\mathbb{E}_{q}[\mathbf{h}_{j,a}]$, which is mitigated by using the previous iteration’s $\mathbb{E}_{q}[u]$ in the calculations, and in the case of it being the first iteration, it is randomly initiated.
III-2 Maximization Step
This step involves determining the estimate of $\mathbf{x}_{j}$ by maximizing $\mathcal{F}(y_{j,a};q,\boldsymbol{\theta})$ with regards to $\boldsymbol{\theta}$, and it is shown in [zhang-deep-2022] that the transmitted signal from AP $j$, may be determined using
$\widehat{\mathbf{x}}_{j}=\operatorname*{argmin}_{\mathbf{x}_{j}}\left[(y_{j,a}-\mathbb{E}_{q}[\mathbf{h}_{j,a}]^{H}\mathbf{x}_{j})^{2}+M\mathbf{x}_{j}^{\,H}\boldsymbol{\Omega}_{\mathbf{h}}\mathbf{x}_{j}\right].$ (11)
These steps are then repeated until the difference of two consecutive iterations of the estimate $\widehat{\mathbf{x}}_{j}$ are sufficiently small, or when the predetermined maximum number of iterations has been achieved.
IV Privacy-Preserving Framework
For a cell-free MIMO ISAC system, we propose a joint, two block, precoder and AP selection framework for privacy-preservation, which utilizes alternating optimization of the two blocks. For a given receiver/transmitter AP configuration, the first block finds the precoding vectors that maximize the SINR at the sensing receivers. The second block minimizes the MI of the received signal at the users, and the sensing signal, by altering the AP configuration. Thereafter, a new set of precoders is generated and the AP configuration is optimized w.r.t. to the new precoder. As a convergence criteria, the procedure stops when two iterations leave the AP configuration unchanged. The first iteration assumes all APs as transmitters.
IV-A Precoder Design
Given a set of APs with predetermined receiver/transmitter duties, the problem of maximization of the received SINR at sensing APs can be formulated as
$\max_{\mathbf{W}}\quad$ $\gamma_{s}(\mathbf{W})$ (12a)
s.t. $\gamma_{\text{UE}_{i}}\geq\gamma,\forall i$ (12b)
$P_{j,p}\leq P_{max},\forall j,\forall p,$ (12c)
where $\mathbf{W}$ is a $MN_{\text{Tx}}\times(N_{\text{UE}}+1)$ matrix consisting of each transmitter APs $M\times(N_{\text{UE}}+1)$ precoder matrix, denoted $\mathbf{W}_{j}$ for AP $j=1,2,...,N_{\text{Tx}}$ [demirhan_cell-free_2023]. In turn, $\mathbf{W}_{j}$ consists of AP $j$’s concatenated precoder vectors for each user plus the sensing precoder, that is $\mathbf{W}_{j}=\left[\mathbf{w}_{j,1},\mathbf{w}_{j,2},...,\mathbf{w}_{j,N_{\text{UE}}},\mathbf{w}_{j,t}\right]$. Hence, the aggregate precoder matrix $\mathbf{W}=\left[\mathbf{W}_{1}^{T},\mathbf{W}_{2}^{T},...,\mathbf{W}_{N_{\text{Tx}}}^{T}\right]^{T}$ columns represent the total precoding vector for a user from every transmitter AP, where the last column represents the target. Further, $\gamma_{\text{UE}_{i}}$ and $\gamma$ represents the total received SINR per user and the minimum SINR threshold for a user, respectively. $P_{j,p}$ is the transmit power from AP $j$’s antenna $p$, and $P_{max}$ is the maximum allowed transmission power for each antenna. Therefore, $P_{j,p}$ can be computed by taking the norm square of each row of $\mathbf{W}$.
Similar to [behdad-multi-static-2023], by taking the norm square of the channel gain of the desired signal in (3), $\gamma_{s}(\mathbf{W})$ may be expanded as
$\gamma_{s}(\mathbf{W})=\frac{\displaystyle\sum\limits_{r=1}^{N_{\text{Rx}}}\sum\limits_{n=1}^{N}\mathbf{s}[n]^{H}\mathbf{W}^{H}\mathbf{A}_{r}\mathbf{W}\mathbf{s}[n]}{\displaystyle N_{\text{Rx}}NM\sigma^{2}_{n}}.$ (13)
by using the the aggregated $MN_{\text{Tx}}\times MN_{\text{Tx}}$ matrix for each receiver $r$
$\mathbf{A}_{r}=\begin{bmatrix}\mathbf{A}_{1,1}&\dots&\mathbf{A}_{1,N_{\text{Tx}}}\\ \vdots&\ddots&\vdots\\ \mathbf{A}_{N_{\text{Tx}},1}&\dots&\mathbf{A}_{N_{\text{Tx}},N_{\text{Tx}}}\end{bmatrix},$
with $\mathbf{A}_{j,m}$ being the $M\times M$ matrices of transmitter AP $j$ and $m$ defined by
$\mathbf{A}_{j,m}=\sqrt{\beta_{j,r}\beta_{m,r}}\mathbf{a}(\theta_{j})^{*}\mathbf{a}(\theta_{r})^{H}\alpha_{j,r}^{*}\alpha_{m,r}\mathbf{a}(\theta_{r})\mathbf{a}(\theta_{m})^{T}.$
In (13), $N$ denotes the number of symbols that are being transmitted to each user, and, consequently, the number of samples in the sensing probing signal. Moreover, $\mathbf{s}[n]$ is the vector of transmitted symbols for all users and the sensing signal, i.e $\left[s_{1}[n],...,s_{N_{\text{UE}}}[n],s_{s}[n]\right]^{T}$.
Since the aim is to optimize $\gamma_{s}$ with respect to $\mathbf{W}$, the Concave-Convex Procedure (CCP) is considered and thus, the objective function must be concave [demirhan_cell-free_2023]. For this purpose the first order Taylor approximation is applied to the term $\mathbf{W}^{H}\mathbf{A}_{r}\mathbf{W}$ around the result of the previous iteration of the procedure, resulting in
$P_{1}^{r}(\mathbf{W}_{p})=2\mathbf{W}^{H}_{p-1}\mathbf{A}_{r}\left(\mathbf{W}_{p}-\mathbf{W}_{p-1}\right)+\mathbf{W}^{H}_{p-1}\mathbf{A}_{r}\mathbf{W}_{p-1}.$ (14)
and thus a final expression for the objective function is obtained as
$\gamma_{s}(\mathbf{W})=\frac{\displaystyle\sum\limits_{r=1}^{N_{\text{Rx}}}\sum\limits_{n=1}^{N}\mathbf{s}[n]^{H}P_{1}^{r}(\mathbf{W})\mathbf{s}[n]}{\displaystyle N_{\text{Rx}}NM\sigma^{2}_{n}}.$ (15)
In (14), $p$ refers to the iteration of the CCP, hence $p-1$ is the $\mathbf{W}$ from the previous iteration and a random matrix is generated for $p=0$.
A similar approach is used to transform the communication SINR constraint into a convex function where the function $\gamma_{\text{UE}_{i}}$ becomes
$\gamma_{\text{UE}_{i}}=\frac{\displaystyle\left|\mathbf{h}_{i}^{H}\mathbf{w}_{i}\right|^{2}}{\displaystyle\left|\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{i}^{H}\mathbf{w}_{k}\right|^{2}+\left|\mathbf{h}_{i}^{H}\mathbf{w}_{t}\right|^{2}+\sigma^{2}_{n}},$ (16)
by using the substitutions $\mathbf{h}_{i}=\left[\mathbf{h}^{T}_{1,i},\mathbf{h}^{T}_{2,i},...,\mathbf{h}^{T}_{N_{\text{Tx}},i}\right]^{T}$ and
$\mathbf{w}_{i}=\left[\mathbf{w}^{T}_{1,i},\mathbf{w}^{T}_{2,i},...,\mathbf{w}^{T}_{N_{\text{Tx}},i}\right]^{T}$, representing the concatenated channel vector and precoder vector to user $i$ from all the transmitting APs, respectively. Again, the first order Taylor approximation is applied around the previous iteration’s value of $\mathbf{w}_{i}$. Also, the slack variables $\tau_{d}$ and $\tau_{n}$ for the denominator and numerator in (16), respectively, are introduced. The final expression for the constraint in (12b) is therefore
$(\mathbf{w}_{i}^{H})_{(p-1)}\mathbf{h}_{i}\mathbf{h}_{i}^{H}(\mathbf{w}_{i})_{(p-1)}\\ +2(\mathbf{w}_{i}^{H})_{(p-1)}\mathbf{h}_{i}\mathbf{h}_{i}^{H}((\mathbf{w}_{i})_{p}-(\mathbf{w}_{i})_{(p-1)})\geq\tau_{n}$ (17a)
$\left|\sum\limits_{\begin{subarray}{c}k=1\\ k\neq i\end{subarray}}^{N_{\text{UE}}}\mathbf{h}_{i}^{H}\mathbf{w}_{k}\right|^{2}+\left|\mathbf{h}_{i}^{H}\mathbf{w}_{t}\right|^{2}+\sigma^{2}_{n}\leq\tau_{d}$ (17b)
$\tau_{n}\geq\tau_{d}\gamma.$ (17c)
IV-B Transmitter and Receiver Access Point Selection
The proposed method to counteract the adversarial user, involves a minimization of the MI w.r.t. APs being ISAC-signal receivers or transmitters. Consider the set of all APs (considered as transmitters), $\mathcal{A}=$$\{a_{l}\}_{1}^{N_{\text{AP}}}$ where $a_{l}$ is each AP $l=1,2,...,N_{AP}$, and $N_{AP}$ is the total number of APs. The problem formulation then becomes
$\min_{\mathcal{R}\subset\mathcal{A}}\quad\sum\limits_{i=1}^{N_{\text{UE}}}\sum\limits_{l=1}^{N_{\text{AP}}}I(y_{i};\mathbf{x}^{s}_{l}),$ (18)
where $\mathcal{R}$ is the set of receiver APs, and $\mathbf{x}^{s}_{l}\in\mathbb{C}^{M\times N}$ is the precoded sensing symbol from each AP $l$, i.e. $\mathbf{x}^{s}_{l}=\mathbf{w}_{l,t}\,\mathbf{s}_{s}^{H}$, for $\mathbf{s}_{s}$ being an $N\times 1$ vector of sensing symbols.
Following [marzetta-fundamentals-2016], the MI can be upper bounded by
$\begin{split}I(y_{i};\mathbf{x}^{s}_{l})&\leq\log_{2}\left(1+\mathbf{h}_{l,i}^{H}\mathbf{x}^{s}_{l}(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right)\\ &=\log_{2}\left(1+\left|\left|(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right|\right|^{2}\right),\end{split}$ (19)
where $\mathbf{h}_{l,i}\in\mathbb{C}^{M}$ is the channel vector between AP $l$ and user $i$. Exploiting the monotonic properties of the $\log_{2}$ function, in conjunction with (19), enables the minimization problem (18) to be rewritten as
$\min_{\mathcal{R}\subset\mathcal{A}}\quad\sum\limits_{i=1}^{N_{\text{UE}}}\sum\limits_{l=1}^{N_{\text{AP}}}\left|\left|(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right|\right|^{2}.$ (20)
Finally, by defining $\mathbf{l}_{\text{Rx}}=\left[l_{1},l_{2},...,l_{N_{\text{Rx}}}\right]^{T}$ as the unique indexes of the receiver APs in $\mathcal{A}$, the solution to (18) becomes
=$\mathcal{R}=\left\{a_{l}\in\mathcal{A}\,\,\forall l\in\mathbf{l}_{\text{Rx}}:\operatorname*{argmax}_{\mathbf{l}_{\text{Rx}}}\sum\limits_{l\in\mathbf{l}_{\text{Rx}}}\sum\limits_{i=1}^{N_{\text{UE}}}\left|\left|(\mathbf{x}^{s}_{l})^{H}\mathbf{h}_{l,i}\right|\right|^{2}\right\}.$ (21)
Equivalently, defining a matrix whose columns are the simplified MI contributions from one AP to all users (and thus the rows are the MI contributions from all APs to one user) as
$\mathbf{M}=\begin{bmatrix}\left|\left|(\mathbf{x}^{s}_{1})^{H}\mathbf{h}_{1,1}\right|\right|^{2}&\dots&\left|\left|(\mathbf{x}^{s}_{N_{\text{AP}}})^{H}\mathbf{h}_{N_{\text{AP}},1}\right|\right|^{2}\\ \vdots&\ddots&\vdots\\ \left|\left|(\mathbf{x}^{s}_{1})^{H}\mathbf{h}_{1,N_{\text{UE}}}\right|\right|^{2}&\dots&\left|\left|(\mathbf{x}^{s}_{N_{\text{AP}}})^{H}\mathbf{h}_{N_{\text{AP}},N_{\text{UE}}}\right|\right|^{2}\end{bmatrix},$
facilitates a modification of the solution into
=$\mathcal{R}=\left\{a_{l}\in\mathcal{A}\,\,\forall l\in\mathbf{l}_{\text{Rx}}:\operatorname*{argmax}_{\mathbf{l}_{\text{Rx}}}\sum\limits_{l\in\mathbf{l}_{\text{Rx}}}\left|\left|\mathbf{M}_{l}\right|\right|\right\},$ (22)
with $\mathbf{M}_{l}$ being the $l$th column vector of $\mathbf{M}$.
Algorithm 1 summarizes the proposed privacy-preserving framework.
Algorithm 1 Privacy-Preserving Framework
0:  Since the framework is proposed for a centralized control system, general system knowledge is assumed.
1:  $\mathcal{R}_{p=0}\leftarrow\emptyset$
2:  $p\leftarrow 0$
3:  repeat
4:     $p\leftarrow p+1$
5:     Solve (12) for $\mathbf{W}_{p}$ by using (15) and (17), for AP receiver configuration $\mathcal{R}_{p-1}$
6:     Generate sensing signals $\mathbf{x}^{s}_{l}$ for each AP $l$, using precoder matrix $\mathbf{W}_{p}$
7:     Get $\mathcal{R}_{p}$ from (22) by using $\mathbf{x}^{s}_{l}$ and $\mathbf{h}_{l,i}$, assuming perfect channel estimation by the centralized control unit
8:  until $\mathcal{R}_{p}=\mathcal{R}_{p-1}$ or $p\geq 50$
8:  $\mathbf{W}_{p},\mathcal{R}_{p}$
V Results
Results presented in this section are produced [^1] https://github.com/ISAC-group-LiU/Privacy-preservation using the parameters in Table I unless specified. The system is randomized, with certain constraints, for each of the $1000$ trials used for each data point in all plots. The effectiveness of the attack model is measured in probability of detection, $P_{D}$, of the target, by the attacker, and is the amount of successful target position estimations, divided by the total number of trials.
TABLE I: Simulation parameters.
Parameter Value Parameter Value
Grid Area $1000\times 1000$ $N_{\text{UE}}$ $3$
Modulation Order $16$ $K$ $4$
Path Loss Exponent $3$ $N$ $16$
Max EM Iterations $100$ $P_{max}$ $35$ dBm
Max CCP Iterations $100$ $\gamma$ $3$ dB
EM Convergence Threshold $1\cdot 10^{-10}$ $\sigma_{H}$ $-40$ dB
CCP Convergence Threshold $0.1$ $\alpha_{j,r}$ $10$ dB
$N_{\text{AP}}$ $6$ $\sigma_{n}$ $-94$ dBm
$N_{\text{Rx}}$ $1$ $\eta$ $1$
$M$ $64$ $R_{\text{GD}}$ $10$
Figure 2: Probability of target detection, with varying min power constraint.
Figure 3: Probability of detection, and sensing SINR, for varying number of receiver APs.
Figure 2 shows $P_{D}$ when using no MI minimization, versus using the framework proposed in this paper, when adjusting the transmit power constraint (12c). As expected, $P_{D}$ is greater when more power is available to the system, due to the sensing beams being more prevalent within the transmitter beampatterns. Moreover, when the framework operates within a system with an abundance of power usage, it performs worse than if the power supply is limited. Again, this is expected considering the sensing beams becoming more difficult to obfuscate when they are stronger.
As seen in Figure 3, $P_{D}$ decreases with more receiver APs, as there are less ISAC transmitters to triangulate an estimated target position. Furthermore, the difference between the two cases increases as more APs are acting as receivers, which demonstrates a privacy-preservation gain from having more degrees of freedom with the developed method. It is worthwhile to mention that, in both cases, the communication SINR and maximum power consumption constraints are met.
Additionally, the effects of the proposed framework on the sensing SINR is shown in Figure 3. The improvement in $P_{D}$ is attained with minimal (to none) sensing SINR loss. Although, the dynamic change on sensing nodes may require increased complexity in the coordination and fusion of sensing information.
VI Conclusions
This paper introduced an attack model in the design of ISAC systems in cell-free MIMO networks. The attacker, a user of the network, leaks sensing information contained in the design of the precoders for ISAC operations. As a countermeasure, we proposed a framework to reduce the probability of detection of the target, within an ISAC system, from the attacker by jointly designing the precoders and selection of sensing units based on mutual information. Our results show that the probability of detection is effectively reduced without significant drawbacks in terms of sensing performance. It is proved that even a simple MI minimization approach can contribute to the prevention of sensing information leakage. While more sophisticated methods can be further analyzed for privacy-preservation on ISAC systems, this paper serves as a benchmark for further development of countermeasures of this type of privacy intrusion.
[…]